# Contrastive Learning Using Graph Embeddings for Domain Adaptation of Language Models in the Process Industry

### v1

@@ Abstract
Recent trends in NLP utilize knowledge graphs (KGs) to enhance pretrained language models by incorporating additional knowledge from the graph structures to learn domain-specific terminology or relationships between documents that might otherwise be overlooked. This paper explores how SciNCL, a graph-aware neighborhood contrastive learning methodology originally designed for scientific publications, can be applied to the process industry domain, where text logs contain crucial information about daily operations and are often structured as sparse KGs. Our experiments demonstrate that language models fine-tuned with triplets derived from GE outperform a state-of-the-art mE5-large text encoder by 9.8-14.3% (5.4-8.0p) on the proprietary process industry text embedding benchmark (PITEB) while being 3-5 times smaller in size.
@@
Anastasia Zhukova1thanks: Equal contribution., Jonas Lührs [^1] footnotemark: 1, Christian E. Matt2, Bela Gipp1 1University of Göttingen, Göttingen, Germany 2eschbach GmbH, Bad Säckingen, Germany anastasia.zhukova@uni-goettingen.de, jonas.luehrs@stud.uni-goettingen.de, christian.matt@eschbach.com, gipp@uni-goettingen.de
@@ 1 Introduction
For several years, domain adaptation of language models (LMs) has followed three state-of-the-art methods: fine-tuning pre-trained LMs using labeled data of domain-specific tasks Devlin et al. (2019); Liu et al. (2019), training domain-specific LMs from scratch Lee et al. (2019); Beltagy et al. (2019); Huang et al. (2019), or domain-adaptive continual pretraining (DAPT) using smaller amounts of domain text data when labeled domain task-specific data is limited Gururangan et al. (2020); Strubell et al. (2020). While DAPT and the subsequent fine-tuning have become a practical and efficient solution for creating specialized NLP applications Guo and Yu (2022), it relies on the availability of unlabeled domain-specific data, from which an LM can learn domain-specific terminology.
@@ 3 Methodology
where $d$ is a document vector representation $f_{lm}(d)=\mathbf{d}$ using an LM, $\xi$ represents the margin that ensures that $\mathbf{d}^{+}$ is at least $\xi$ closer to $\mathbf{d}^{\mathcal{Q}}$ than $\mathbf{d}^{-}$, and $\|\Delta d\|_{2}$ is the $L^{2}$ norm, which is used as a distance function. In our experiments, we use $d$ as a vector of [CLS] token from the last hidden layer or a concatenation of [CLS] and mean pooling of all tokens from the last hidden layer.
@@ 4.1 Experiment setup
For LM fine-tuning, we selected several pre-trained and domain-adapted mono-lingual or multilingual pre-trained LMs. Specifically, we selected state-of-the-art LMs with BERT architectures that support the German language: GBERT-base Chan et al. (2020), multilingual mBERT-base Devlin et al. (2019), and XLM-RoBERTa-base Conneau et al. (2020). Additionally, we evaluate a domain LM tailored for the process industry using continual pretraining daGBERT Zhukova et al. (2025b). All selected models have an embedding size of 768, and the number of parameters is under 300M.
@@
We evaluate our models and baselines using a private, manually verified process industry text embedding benchmark (PITEB) Zhukova et al. (2025a). Table˜2 provides details on the test collection, which comprises data from seven production plants in the chemical and pharmaceutical domains, totaling 205 queries across a collection of 330K documents. Data from four plants is used to build training datasets for both document-level triplets and a dataset for bi-encoders.
@@
Table 2: Plant industry text embedding test collection (PITEB). The data from three plants was unseen during the LM fine-tuning on the document level.
@@
ibm-granite/granite-embedding-107m-multilingual 107 -
@@
ibm-granite/granite-embedding-278m-multilingual 278 -
@@
We evaluate the text encoder with three metrics: mean average precision (MAP@10), mean reciprocal rank (MRR), and normalized discounted cumulative gain (nDCG@10). Using all three metrics together provides a more well-rounded evaluation, giving insights into ranking quality, early retrieval performance, and overall retrieval quality. We report the mean of these metrics across seven plants.
@@ 4.2 Results
shows the breakdown of performance at the plant level. While some text encoder baselines achieve the best performance on the data from individual plants, no baseline system consistently dominates across all plants in the PITEB. In contrast, the models fine-tuned for the process industry domain generally perform systematically better across all plants. Notably, we observe a positive trend in plants whose data was not used for training. However, despite these improvements, further investigation is necessary to refine the methodology and achieve optimal performance for each plant.
@@ Acknowledgments
This project is supported by the Federal Ministry for Economic Affairs and Climate Action (BMWK) on the basis of a decision by the German Bundestag. Additional funding was provided by the Federal Ministry of Education and Research (BMBF) in the form of a fellowship under the IFI program of the German Academic Exchange Service (DAAD). We would like to thank Marco Kaiser and Thomas Walton for their contributions to the link prediction services. We also appreciate Malte Ostendorff for the insightful discussions on solutions for domain adaptation in SciNCL.
@@ A.1 Knowledge graph
Data analysis revealed that many FLs are referenced in text logs using domain-specific abbreviations (e.g., A11) but are not explicitly linked as attributes (see Figure˜1). This lack of direct association limits our implementation’s ability to construct a comprehensive domain KG. To address this problem, we implement two domain-specific custom link prediction models to (1) reconstruct the reports_about and related_to connections within the domain KG, (2) expand the context of domain-specific abbreviations. In the context expansion, we replace the abbreviations such as "A11" with a combination of the abbreviation and its description "A11 Pumpe", which was obtained from the linked FL node. As shown in Table˜5, this approach significantly increases the number of graph edges and expands the number of text log nodes, i.e., text logs that previously lacked associated FLs are now integrated into the graph, enhancing its overall structure and connectivity. Moreover, we apply a record linking model Zhukova et al. (2025c) to restore connectivity between the text logs, i.e., when two records reported about a problem and a solution to it, but were not explicitly linked in a database as related. Table˜5 shows that the KG after the link prediction has the largest number of text log nodes used up from the available raw data.
@@ A.2 Graph embeddings
Our methodology primarily adopts the training parameters for the graph embedding model described in SciNCL, while experimenting with variations in the number of training epochs and using sentence embedding initialization. Some FLs have similar descriptions, such as reactors of different types. Initializing the graph network with semantically meaningful vectors can enhance the learning of structural connections between nodes. For example, by initializing FL nodes using their descriptions, similar FL types (e.g., pumps) are naturally positioned closer together in the embedding space. During training, text logs referencing different pumps are more likely to be mapped to similar regions within the embedding space, as their corresponding FLs were initially placed near each other. Consequently, text logs that do not share the same FLs but originate from semantically similar contexts can still achieve similar node embeddings. Moreover, if text logs contain terms similar to those in the descriptions of FLs, such nodes will also be positioned closely in the vector space. We anticipate that integrating domain-specific relationships with semantic similarity will help offset challenges such as low graph connectivity and poor text quality while improving the sampling of similar and dissimilar text documents.
The node embeddings for the graph model are initialized by vectorizing the text logs and functional location (FL) descriptions using sentence transformer embeddings from PM-AI/bi-encoder_msmarco_bert-base_german. Table˜6 presents the link prediction performance for plant D, evaluated using metrics such as MRR, Hits@1, Hits@10, and AUC. For these experiments, 99% of the data was used for training, while the remaining 1% (1193 edges) was reserved for testing [^2] Metric definitions are available in the PBG documentation: https://torchbiggraph.readthedocs.io/en/latest/evaluation.html. The results demonstrate that initializing nodes with text embeddings significantly enhances link prediction performance. These findings underscore the critical role of initializing node embeddings with text embeddings for effectively training graph embedding models in the process industry.
@@ A.3 Triplet sampling
To identify text logs with a similar graph context — those associated with similar text logs and functional locations — an approximate nearest neighbor (ANN) search was conducted in the text log embedding space. FAISS Johnson et al. (2021) was employed to construct a flat index of text logs, keyed by their node embeddings. To ensure consistency in distance measurement, embeddings were L2-normalized, and cosine similarity was used as the distance metric by computing the inner product. After building and populating the FAISS index with text log embeddings, queries in the form of text log embeddings $\mathbf{d}^{Q}$ were used to retrieve the $k$-nearest neighbors based on their cosine distance in the embedding space.
@@
Sampling strategies differ between positive and types of negative samples, i.e., similar to SciNCL, we use kNN sampling for positives and hard negatives, and filtered random sampling for easy negatives. kNN sampling is conducted as k-nearest neighbors search $kNN(f_{gem}(\mathbf{d}^{Q}),\mathbf{D})$ given a graph embedding model (e.g., PyTorch BigGraph Lerer et al. (2019)) denoted as $f_{gem}$, document node embeddings $\mathbf{D}$, and a search index (e.g., FAISS Johnson et al. (2021)). From the neighbors around the query document $d^{Q}$, $c^{\prime}$ samples are selected using the interval $(k-c^{\prime},k]$, where $N=\{n_{1},n_{2},n_{3},\dots\}$ represents the neighbors, i.e., $n_{i}$ is the $i$-th nearest neighbor in the graph embedding space. For example, if $c^{\prime}=3$ and $k=10$, the selected samples would be the three closest neighbors within the specified range: $n_{8},n_{9}$, and $n_{10}$. Filtered random sampling utilizes random sampling, i.e., samples $c^{\prime}$ documents from the corpus without replacement, but excludes the documents retrieved by kNN, i.e., neighbors within the largest $k$.
Due to the low number of direct connections between text logs, we select the closest $k=2$ neighbors as positive samples. A small $k^{+}$ ensures that text logs which are positioned close in the graph embedding space due to their direct links are included. A higher $k^{+}$ value could risk skipping such direct links. Similarly, $k_{hard}^{-}$ was set to 50 to ensure that the selected hard negatives were sufficiently dissimilar, as the 50th neighbor is expected to fall outside the direct neighborhood of the query log.
@@ A.4 Domain-related MS MARCO Dataset
We created a domain-related version of MS MARCO by running the document collection through our binary classifier that identifies if a document is domain-related or not. This classifier is a fine-tuned SciBERT model Beltagy et al. (2019) trained on an 80K dataset that combines fields of study (FoS) data and our in-domain data. The train dataset was split so that 50% consisted of domain-related texts, selected by the labels in FoS that correspond to our domain, and in-domain texts, thereby forming the true labels, while the other 50% was sampled from various areas of FoS. We classified the English documents and used the document IDs to extract the German version of the same texts. A document was considered domain-related if a confidence in the positive label was larger than 0.9, and we collected the query-document pairs where the documents were labeled as domain-related. The resulting dataset contained approximately 2.27M training pairs, where 132K pairs were positive.

### v2

@@ Abstract
Recent trends in NLP utilize knowledge graphs (KGs) to enhance pretrained language models by incorporating additional knowledge from the graph structures to learn domain-specific terminology or relationships between documents that might otherwise be overlooked. This paper explores how SciNCL, a graph-aware neighborhood contrastive learning methodology originally designed for scientific publications, can be applied to the process industry domain, where text logs contain crucial information about daily operations and are often structured as sparse KGs. Our experiments demonstrate that language models fine-tuned with triplets derived from graph embeddings (GE) outperform a state-of-the-art mE5-large text encoder by 9.8-14.3% (5.45-7.96p) on the proprietary process industry text embedding benchmark (PITEB) while having 3 times fewer parameters.
@@
Anastasia Zhukova1thanks: Equal contribution., Jonas Lührs [^1] footnotemark: 1, Christian E. Lobmüller2, Bela Gipp1 1University of Göttingen, Göttingen, Germany 2eschbach GmbH, Bad Säckingen, Germany anastasia.zhukova@uni-goettingen.de, jonas.luehrs@stud.uni-goettingen.de, christian.matt@eschbach.com, gipp@uni-goettingen.de
@@ 1 Introduction
For several years, domain adaptation of language models (LMs) has followed three well-established approaches: fine-tuning pre-trained LMs using labeled data of domain-specific tasks Devlin et al. (2019); Liu et al. (2019), training domain-specific LMs from scratch Lee et al. (2019); Beltagy et al. (2019); Huang et al. (2019), or domain-adaptive continual pretraining (DAPT) using smaller amounts of domain text data when labeled domain task-specific data is limited Gururangan et al. (2020); Strubell et al. (2020). While DAPT and the subsequent fine-tuning have become a practical and efficient solution for creating specialized NLP applications Guo and Yu (2022), it relies on the availability of unlabeled domain-specific data, from which an LM can learn domain-specific terminology.
@@ 3 Methodology
where $d$ is a document vector representation $f_{lm}(d)=\mathbf{d}$ using an LM, $\xi$ represents the margin that ensures that $\mathbf{d}^{+}$ is at least $\xi$ closer to $\mathbf{d}^{\mathcal{Q}}$ than $\mathbf{d}^{-}$, and $\|\Delta d\|_{2}$ is the $L^{2}$ norm, which is used as a distance function. In our experiments, we use $d$ as (1) a vector of [CLS] token from the last hidden layer or (2) a concatenation of [CLS] and mean pooling of all tokens from the last hidden layer.
@@ 4.1 Experiment setup
For LM fine-tuning, we selected several pre-trained and domain-adapted mono-lingual or multilingual pre-trained LMs. Specifically, we selected state-of-the-art LMs with BERT architectures that support the German language: GBERT-base Chan et al. (2020), multilingual mBERT-base Devlin et al. (2019), and XLM-RoBERTa-base Conneau et al. (2020). Additionally, we evaluate daGBERT, a domain LM tailored for the process industry using continual pretraining Zhukova et al. (2025b). All selected models have an embedding size of 768, and the number of parameters is under 300M.
@@
We evaluate our models and baselines using a private expert-verified process industry text embedding benchmark (PITEB) Zhukova et al. (2025a). Table˜2 provides details on the test collection, which comprises data from seven production plants in the chemical and pharmaceutical domains, totaling 205 queries across a collection of 330K documents. Data from four plants is used to build training datasets for both document-level triplets and a dataset for bi-encoders.
@@
Table 2: Process industry text embedding benchmark (PITEB) for semantic search evaluation. The data from three plants was unseen during the LM fine-tuning on the document level.
@@
ibm-granite/granite-embedding-107m-multiling. 107 -
@@
ibm-granite/granite-embedding-278m-multiling. 278 -
@@
We evaluate the text encoder with three metrics: mean average precision (MAP@10), mean reciprocal rank (MRR@10), and normalized discounted cumulative gain (nDCG@10). Using all three metrics together provides a more well-rounded evaluation, giving insights into ranking quality, early retrieval performance, and overall retrieval quality. We report the mean of these metrics across seven plants.
@@ Acknowledgments
This project is supported by the Federal Ministry for Economic Affairs and Climate Action (BMWK) on the basis of a decision by the German Bundestag. Additional funding was provided by the Federal Ministry of Education and Research (BMBF) in the form of a fellowship under the IFI program of the German Academic Exchange Service (DAAD). We would like to thank Marco Kaiser and Thomas Walton for their contributions to the link prediction services. We also appreciate Malte Ostendorff for the insightful discussions on solutions towards domain adaptation using SciNCL.
@@ A.1 Knowledge graph
Data analysis revealed that many FLs are referenced in text logs using domain-specific abbreviations (e.g., A11) but are not explicitly linked as attributes (see Figure˜1). This lack of direct association limits our implementation’s ability to construct a comprehensive domain KG. To address this problem, we implement two domain-specific custom link prediction models to (1) reconstruct the reports_about and related_to connections within the domain KG, (2) expand the context of domain-specific abbreviations. In the context expansion, we replace the abbreviations such as "A11" with a combination of the abbreviation and its description "A11 Pumpe", which was obtained from the linked FL node. As shown in Table˜5, this approach significantly increases the number of graph edges and expands the number of text log nodes, i.e., text logs that previously lacked associated FLs are now integrated into the graph, enhancing its overall structure and connectivity. Moreover, we apply a record linking model Zhukova et al. (2025c) to restore connectivity between the text logs, i.e., when two records reported about a problem and a solution to it, but were not explicitly linked in a database as related. Table˜5 shows that the KG after the link prediction includes the largest number of text log nodes from the available raw data.
@@ A.2 Graph embeddings
Our methodology primarily adopts the training parameters for the graph embedding model described in SciNCL, while experimenting with variations in the number of training epochs and using sentence embedding initialization. Some FLs have similar descriptions, such as reactors of different types. Initializing the graph network with semantically meaningful vectors can enhance the learning of structural connections between nodes. For example, by initializing FL nodes using their descriptions, similar FL types (e.g., pumps) are naturally positioned closer together in the embedding space. During training, text logs referencing different pumps are more likely to be mapped to similar regions within the embedding space, as their corresponding FLs were initially placed near each other. Consequently, text logs that do not share the same FLs but originate from semantically similar contexts can still achieve similar node embeddings. Moreover, if text logs contain terms similar to those in the descriptions of FLs, such nodes will also be positioned closely in the vector space. We anticipate that integrating domain-specific relationships with semantic similarity will help mitigate challenges such as low graph connectivity and poor text quality, while improving the sampling of both similar and dissimilar text documents.
The node embeddings for the graph model are initialized by vectorizing the text logs and functional location (FL) descriptions using sentence transformer embeddings from PM-AI/bi-encoder_msmarco_bert-base_german. Table˜6 presents the link prediction performance for plant D, evaluated using metrics such as MRR, Hits@1, Hits@10, and AUC. For these experiments, 99% of the data was used for training, while the remaining 1% (1193 edges) was reserved for testing [^2] Metric definitions are available in the PBG documentation: https://torchbiggraph.readthedocs.io/en/latest/evaluation.html. The results demonstrate that initializing nodes with text embeddings significantly enhances the performance of link prediction. These findings highlight the crucial role of initializing node embeddings with text embeddings in effectively training graph embedding models within the process industry.
@@ A.3 Triplet sampling
To identify text logs with a similar graph context — those associated with similar text logs and functional locations — an approximate nearest neighbor (ANN) search was conducted in the text log embedding space. FAISS Johnson et al. (2021) was employed to construct a flat index of text logs, keyed by their node embeddings. To ensure consistency in distance measurement, embeddings were L2-normalized, and cosine similarity was used as the distance metric by computing the inner product of the embeddings. After building and populating the FAISS index with text log embeddings, queries in the form of text log embeddings $\mathbf{d}^{Q}$ were used to retrieve the $k$-nearest neighbors based on their cosine distance in the embedding space.
@@
Sampling strategies differ between positive and types of negative samples, i.e., similar to SciNCL. We use kNN sampling for positives and hard negatives, and filtered random sampling for easy negatives. kNN sampling is conducted as k-nearest neighbors search $kNN(f_{gem}(\mathbf{d}^{Q}),\mathbf{D})$ given a graph embedding model (e.g., PyTorch BigGraph Lerer et al. (2019)) denoted as $f_{gem}$, document node embeddings $\mathbf{D}$, and a search index (e.g., FAISS Johnson et al. (2021)). From the neighbors around the query document $d^{Q}$, $c^{\prime}$ samples are selected using the interval $(k-c^{\prime},k]$, where $N=\{n_{1},n_{2},n_{3},\dots\}$ represents the neighbors, i.e., $n_{i}$ is the $i$-th nearest neighbor in the graph embedding space. For example, if $c^{\prime}=3$ and $k=10$, the selected samples would be the three closest neighbors within the specified range: $n_{8},n_{9}$, and $n_{10}$. Filtered random sampling utilizes random sampling, i.e., samples $c^{\prime}$ documents from the corpus without replacement, but excludes the documents retrieved by kNN, i.e., neighbors within the largest $k$.
Due to the low number of direct connections between text logs, we select the closest $k=2$ neighbors as positive samples. A small $k^{+}$ ensures that text logs that are positioned close in the graph embedding space, due to their direct links, are included. A higher $k^{+}$ value could risk skipping such direct links. Similarly, $k_{hard}^{-}$ was set to 50 to ensure that the selected hard negatives were sufficiently dissimilar, as the 50th neighbor is expected to fall outside the direct neighborhood of the query log.
@@ A.4 Domain-related MS MARCO Dataset
We created a domain-related version of MS MARCO by running the document collection through our binary classifier, which identifies whether a document is domain-related or not. This classifier is a fine-tuned SciBERT model Beltagy et al. (2019) trained on an 80K dataset that combines fields of study (FoS) data and our in-domain data. The train dataset was split so that 50% consisted of domain-related texts, selected by the labels in FoS that correspond to our domain, and in-domain texts, thereby forming the true labels, while the other 50% was sampled from various areas of FoS. We classified the English documents and used the document IDs to extract the German version of the same texts. A document was considered domain-related if a confidence in the positive label was larger than 0.9, and we collected the query-document pairs where the documents were labeled as domain-related. The resulting dataset contained approximately 2.27M training pairs, where 132K pairs were positive.